OpenStudy (anonymous):

Please help i'm stuck...... To produce a particular radio model, it costs a manufacturer $30 per radio, and it is assumed that if 500 radios are produced, all of them will be sold. What must be the selling price per radio to ensure that the profit (revenue from the sales minus the total production cost) on the 500 radios is greater than $8,200?

jimthompson5910 (jim_thompson5910):

where are you stuck at?

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

what is x in this case

OpenStudy (anonymous):

revenue

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

if we wanted the profit to be 8,200 then we would say profit = (revenue) - (cost) 8,200 = x - 15,000 agreed?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

solve `8,200 = x - 15,000` for x to get x = ???

OpenStudy (anonymous):

23200

jimthompson5910 (jim_thompson5910):

so the total revenue must be 23,200 if you want a profit of 8,200

jimthompson5910 (jim_thompson5910):

now use the idea revenue = (number of units sold)*(price per unit)

OpenStudy (mikeymaximum):

Anything that is $47 or more

OpenStudy (anonymous):

23,200= (500)(30)

jimthompson5910 (jim_thompson5910):

`23,200= (500)(30)` is a false equation

jimthompson5910 (jim_thompson5910):

$30 is the cost and NOT the selling price

jimthompson5910 (jim_thompson5910):

let y = selling price revenue = (number of units sold)*(price per unit) 8200 = (500)*(y) y = ???

OpenStudy (anonymous):

46.4

OpenStudy (anonymous):

so anything greater than 46.4

jimthompson5910 (jim_thompson5910):

idk how you got 46.4

jimthompson5910 (jim_thompson5910):

oh sorry I put in the wrong revenue, my bad

jimthompson5910 (jim_thompson5910):

revenue = (number of units sold)*(price per unit) 23200 = (500)*(y) y = 46.4 so you have the correct selling price

OpenStudy (anonymous):

thank u

jimthompson5910 (jim_thompson5910):

no problem